### v1

# Smectic Liquid Crystal in Modified Landau-de Gennes Model

Abstract
We study the behaviour of Smectic-A (SmA) liquid crystals with a modified Landau-de Gennes (mLdG) model. The director and orientational order of SmA are described by a tensor-order parameter $\mathbf{Q}$, and the positional order is described by a real scalar $u$, which represents the deviation from the average concentration of liquid crystal molecules. Firstly, we prove the existence and regularity of the global minimiser of the mLdG free energy. Then, we analytically prove that the mLdG model can exhibit Isotropic-Nematic-Smectic phase transition in terms of the critical temperatures and stabilities of phases. Furthermore, we explore smectic configurations confined within a square domain with edge length $\lambda$. We analytically study the positional order far from and near the defects and the limiting cases with $\lambda\to 0$ and $\lambda\to\infty$, to potentially guide the behavior of SmA in finite domain size. With finite $\lambda$, we numerically recover BD-like and D-like states observed in experiments. The stability of a BD-like state is enhanced by the layer structure of positional order. The analogous $\mathbf{Q}$ profile of D-like states in nematics and smectics indicates the memory of the director in the N-S phase transition. We also study the frustrated smectic energy landscape and the strategy to alleviate the frustration.
1 Introduction
Liquid crystals are mesophases intermediate between the solid and liquid states, characterized by orderly molecular arrangements [1], that is, the molecules tend to align along certain locally preferred directions, referred to as “directors” in the literature. These orderly molecular arrangements give rise to distinctive optical and electrical properties in liquid crystals, making them valuable in display technologies, optical devices, and sensors [2, 3, 4, 5]. Liquid crystals can exhibit different phases, such as nematic and smectic phases. The nematic phase has long-range orientational order but lacks positional order, while the smectic phase possesses both long-range orientational order and positional order, leading to a layered structure with positional coherence within the layers [6]. There are also several smectic phases, such as Smectic-A and Smectic-C, each with distinct characteristics [7]. In the Smectic-A phase, the director is parallel to the normal of the layer. In contrast, in the Smectic-C phase, there is a non-zero angle between the director and the normal of the layer. In this paper, we focus on the Smectic-A phase, which will be simply referred to as the “smectic phase”.
External constraints, such as confinement and boundary anchoring, can induce deformations in the liquid crystal. These deformations may not coincide with the liquid crystal phase in the bulk, leading to geometric frustrations. As a result, a diverse array of textures with characteristic defect structures may spontaneously assemble [8, 9, 10]. For instance, when smectic is deposited on a substrate that promotes varying boundary anchoring, their layers may bend and form focal conic domains (FCDs) [11, 12]. These FCDs have been utilized as guides for colloidal dispersion [13], in soft lithography [14], and as templates for superhydrophobic surfaces [15]. The experimental observations in [16] suggest a stable BD-type smectic profile confined on a square, where the molecules align either horizontally or vertically, producing two line defects near the opposite edges of the square. This BD-type configuration found to be stable in confined smectic phases is considered unstable in nematic [17, 18, 19], which indicates the distinctiveness of the properties of confined smectic configurations.
Recent discoveries in the study of Nematic-Smectic (N-S) phase transition have highlighted the interplay between the director and layer structure. For example, the existence of geometric memory in the N-S transition leads to FCDs melting into a dense array of boojums defects [20]. By using colloidal silica rods and leveraging their significant density difference with the dispersing solvent, nematic and smectic phases can be confined within a single chamber which produces a smectic-nematic interface, and the directors in the smectic-nematic interface impose their profile in the nematic slice [16]. Under shell confinement, the N-S phase transition, or the emergence of the layer structure, initially occurs on the thicker side of the shell, distant from the point defects [21]. These experimental findings inspire us to study the N-S phase transition and the confined smectic configurations.
The very complicated structures that emerge in the frustrated smectic phase are challenging to model mathematically. The key point in modelling the smectic phase is to build the nematic director and the order of layer structure, i.e. an additional positional order parameter must be introduced to describe the modulation of the concentration compared to a more simple nematic phase. In recent decades, several powerful continuum mathematical theories for nematic director from microscopic models to macroscopic models have been proposed, such as the microscopic Onsager model, the macroscopic Landau-de Gennes (LdG) model, the macroscopic Oseen-Frank model, and the Ericksen-Leslie model [22]. For modelling the smectic phase, an additional positional order is required to construct the layered structure. For instance, the extended Maier-Saupe model [23] is a molecular model for the smectic phase, which qualitatively predicts the N-S phase transition as a function of temperature. The molecular model is physical, and the parameters can be associated with the molecular structure, but the molecular model is thought to be computationally challenging due to its inherent high-dimensional complexity. For computational convenience, there have also been some phenomenological frameworks for adding the density modulation in Oseen-Frank energy or Landau-de Gennes energy for a nematic, to model smectic phase [24, 25, 26, 27], which successfully predict the structures observed in experiments. However, most of the existing results in these papers focus on numerical results, with a lack of interpretability of the models. For instance, what is the influence of parameters on the stable states (or global energy minimisers), and can the model depict the N-S phase transition with respect to temperature? To address these questions, we investigate the modified Landau-de Gennes (mLdG) model as presented in [27], which is adept at capturing geometric frustration, FCDs, and oily streaks [28], commonly observed in the experiments of confined smectics.
In this paper, we study the smectic liquid crystal in both phase transition and confinement problems using the mLdG model, which incorporates the order of the director $\mathbf{Q}$ and the positional order $u$. First, we prove the existence and regularity of the minimiser, taking into account Dirichlet tangential boundary conditions and weak tangential anchoring boundary conditions. Subsequently, we analytically study the Isotropic-Nematic-Smectic (I-N-S) phase transitions as a function of temperature, and we prove that the nematic phase becomes increasingly unstable and bifurcates into a more stable smectic phase as the temperature decreases. Leveraging the regularity result, we demonstrate that the coupling term prefers not to have a layer structure near the defect and favours a layered structure when $\mathbf{Q}$ is uniaxial. Last, we study the smectic configuration confined on a square with edge length $\lambda$. As $\lambda\rightarrow 0$, we analytically compute the solution of the Euler-Lagrange equation. This solution reveals a Well Order Reconstruction Solution (WORS)-type director field [29], characterized by either horizontal or vertical alignment of directors, and a linear $u$ which does not have layer structure, indicating that the liquid crystal cannot form into layer structure within a very small domain size. As $\lambda$ tends towards infinity, we provide preliminary evidence that the number of layers will increase to infinity. For finite but non-zero domain sizes, on the one hand, the globally stable smectic configuration exhibits WORS-type, Boundary Distortion (BD)-type [18], and Diagonal (D)-type directors as the domain size increases, which aligns with experimental observations in [16]. On the other hand, the N-S structural transition under square confinement with moderate domain size as temperature decreases is investigated. In addition, we find multiple almost uniaxial stable smectic states and identify the transition pathways between them, which reveals a frustrating energy landscape in the density distribution $u$, suggesting the existence of numerous metastable states with only minor differences in density.
This paper is structured as follows: Section 2 provides a brief overview of the LdG and mLdG models, and we analytically establish the existence and regularity of the global minimiser. Section 3 is devoted to the study of I-N-S thermotropic phase transitions. In Section 4, we explore the smectic configurations in confinement. Finally, we present our conclusion and discussion.
2 Theoretical framework
The Landau-de Gennes (LdG) model [1] is the most celebrated continuum theory for nematic liquid crystals and has been hugely successful for describing the Isotropic-Nematic (I-N) phase transition [30] and structural transitions for nematics [31]. The LdG theory describes the nematic phase by the LdG $\mathbf{Q}$-tensor order parameter, which is a traceless and symmetric $3\times 3$ matrix. The $\mathbf{Q}$ tensor is isotropic if $\mathbf{Q}=0$, uniaxial if $\mathbf{Q}$ has a pair of degenerate nonzero eigenvalues, and biaxial if $\mathbf{Q}$ has three distinct eigenvalues [1]. A uniaxial nematic phase has a single distinguished direction of averaged molecular alignment, modelled by the eigenvector or director with the non-degenerate eigenvalue. A biaxial nematic phase has a primary and secondary nematic director i.e. at least two distinguished material directions. In approximately two-dimensional (2D) scenarios, we can use the reduced Landau-de Gennes (rLdG) model, with the rLdG order parameter - a symmetric and traceless $2\times 2$ matrix with only two degrees of freedom: one degree of freedom for the nematic director in the plane and the second degree of freedom describes the degree of ordering about the 2D director [32, 18, 33, 34]. In this paper, we use a modified LdG (mLdG) theory to study confined smectic phases, wherein we use either the LdG or the rLdG order parameter to describe the orientational/nematic ordering with an additional real-valued positional order parameter $u$ and additional energy terms to manipulate the layer structure of $u$ and the coupling between orientational and positional order [27, 35]. More precisely, $u$ models the deviation from the average concentration of the liquid crystal molecules, and a periodic full cycle of $u$ models a SmA layer.
2.1 Preliminaries
The modified Landau-de Gennes (mLdG) energy [27, 35] is given by
(1) $E(\mathbf{Q},u)=\int_{\Omega}\left(f_{LdG}(\mathbf{Q})+f_{bs}(u)+f_{int}(\mathbf{Q},u)\right)\mathrm{d}\mathbf{x},$
where $\Omega\subset\mathbb{R}^{d}$ ($d=2,3$) is the working domain, the nematic order parameter $\mathbf{Q}(\mathbf{x})\in\mathbb{R}^{d\times d}$, and the positional order $u(\mathbf{x})\in\mathbb{R}$. $\mathbb{R}^{d\times d}$ is the space of $d\times d$ matrices.
The first term in (1) is the LdG free energy density in $d$-dimensions (where $d=2,3$),
(2) $f_{LdG}(\mathbf{Q}):=\frac{K}{2}\left|\nabla\mathbf{Q}\right|^{2}+f_{bn}\left(\mathbf{Q}\right),$
where $K$ is a positive material-dependent elastic constant. The elastic energy density penalizes spatial inhomogeneities, and the thermotropic bulk energy density, $f_{bn}$, dictates the preferred NLC phase as a function of temperature,
(3) $\begin{cases}&f_{bn}(\mathbf{Q}):=\frac{A}{2}\mathrm{tr}\mathbf{Q}^{2}-\frac{B}{3}\mathrm{tr}\mathbf{Q}^{3}+\frac{C}{4}(\mathrm{tr}\mathbf{Q}^{2})^{2},d=3,\\ &f_{bn}(\mathbf{Q}):=\frac{A}{2}\mathrm{tr}\mathbf{Q}^{2}+\frac{C}{4}(\mathrm{tr}\mathbf{Q}^{2})^{2},d=2,\end{cases}$
where $A=\alpha_{1}(T-T_{1}^{*})$ is the rescaled temperature, with $\alpha_{1}>0$ and $T_{1}^{*}$ is a characteristic liquid crystal temperature; $B,C>0$ are material-dependent bulk constants. For example, typical values for the representative NLC material MBBA are $B=0.64\times 10^{4}\text{Nm}^{-2}$, $C=0.35\times 10^{4}\text{Nm}^{-2}$ and $K=4\times 10^{-11}\text{N}$ [31, 36]. The minimisers of $f_{bn}$ depend on $A$ and determine the NLC phase for spatially homogeneous samples. In three-dimensions (3D), the minimiser of $f_{bn}$ is the isotropic state for $A>\frac{B^{2}}{27C}$ and in the 2D case, the minimiser of $f_{bn}$ is the isotropic state for $A>0$. For $A<0$, the minimisers of $f_{bn}$ constitute a continuum of $\mathbf{Q}$-tensors defined below:
$\mathcal{N}=\begin{cases}\left\{\mathbf{Q}=s_{+}\left(\mathbf{n}\otimes\mathbf{n}-\frac{\mathbf{I}}{3}\right)\right\},d=3,\\ \left\{\mathbf{Q}=s_{+}\left(\mathbf{n}\otimes\mathbf{n}-\frac{\mathbf{I}}{2}\right)\right\},d=2,\end{cases}$
where
$s_{+}=\begin{cases}\frac{B+\sqrt{B^{2}-24AC}}{4C},A<\frac{B^{2}}{27C},d=3,\\ \sqrt{\frac{-2A}{C}},A<0,d=2,\end{cases}$
and $\mathbf{n}$ is an arbitrary unit vector field (referred to as the nematic director), and $\mathbf{I}_{d}$ is the $d\times d$ identity matrix.
The second term in (1) is the bulk energy density of the smectic order parameter $u$:
(4) $f_{bs}(u)=\frac{a}{2}u^{2}+\frac{b}{3}u^{3}+\frac{c}{4}u^{4},$
where $a=\alpha_{2}(T-T^{*}_{2})$ is a temperature-dependent parameter with $\alpha_{2}>0$, and $T_{2}^{*}<T_{1}^{*}$ is a critical material temperature related to N-S phase transition; $b,c>0$ are material-dependent constants. A non-zero $b$ will result in a non-symmetrical layer structures [25], and we take $b=0$ to study symmetric layer structures. When $a<0$, i.e. the temperature is low enough, the minimisers of $f_{bs}(u)$ prefer a non-zero density distribution $u$.
The third term in (1) is the coupling term between the smectic and nematic order parameters:,
(5) $f_{int}(\mathbf{Q},u)=\begin{cases}B_{0}\left|D^{2}u\right|^{2},A\geqslant\frac{B^{2}}{27C},d=3\text{ }or\text{ }A\geqslant 0,d=2,\\ B_{0}\left|D^{2}u+q^{2}\left(\frac{\mathbf{Q}}{s_{+}}+\frac{\mathbf{I}_{d}}{d}\right)u\right|^{2},otherwise,\end{cases}$
where $B_{0}$ is a phenomenological coupling constant between $\mathbf{Q}$ and $u$, $D^{2}u$ is the Hessian of $u$ and $q$ is a phenomenological parameter [25, 27]. In subsequent sections, we use heuristic arguments to show that for energy minimisers, we expect $q$ to be related to the SmA layer thickness $d$ by $q=\frac{2\pi}{d}$, and hence, $q$ is often identified with the wave number of the SmA layers [25, 27]. The layer thickness $d$ of a homogenous SmA, is usually slightly larger than the long axis of a rod-like liquid crystal molecule, $L$, but less than $2L$ [37]. The layer thickness of the equal mass mixture of 8OPhPy8 and 6OPhPy8 in the SmA phase is about $28.5$ Angstrom in [38].
We consider the full LdG model in the 3D case, in this section, to prove generic existence and regularity results for minimisers of the mLdG energy. Analogous results hold for the rLdG model with minor adaptations of the arguments. The admissible $\mathbf{Q}$-tensors belong to the space
(6) $W^{1,2}_{\Omega,\mathbf{S}_{0}}=\{\mathbf{Q}\in\mathbf{S}_{0}|\mathbf{Q}\in W_{\Omega}^{1,2}\},$
and the admissible smectic order parameter, $u$, belongs to $W^{2,2}_{\Omega}$, where
(7) $\mathbf{S}_{0}:=$$\{\mathbf{Q}\in\mathbb{R}^{3\times 3}:\mathbf{Q}_{ij}=\mathbf{Q}_{ji},\sum_{i=1}^{3}\mathbf{Q}_{ii}=0\},$
$W_{\Omega}^{k,p}=$$\left\{u:\int_{\Omega}\left(|u|^{p}+\sum_{|\alpha|\leqslant k}|D^{\alpha}u|^{p}\right)\mathrm{d}x<\infty\right\}.$
To study the Isotropic-Nematic-Smectic phase transition and structural transitions for SmA in confinement, we consider three different kinds of boundary conditions: (1) Periodic boundary condition for $\mathbf{Q}$ and $u$ on a one-dimensional domain $\Omega=[0,h]$:
(8) $\begin{cases}\mathbf{Q}(0)=\mathbf{Q}(h),D_{x}\mathbf{Q}(0)=D_{x}\mathbf{Q}(h),\\ u(0)=u(h),D_{x}u(0)=D_{x}u(h),D_{xx}u(0)=D_{xx}u(h).\end{cases}$
We impose periodic boundary conditions on the derivative of $\mathbf{Q}$ to ensure that $\mathbf{Q}$ is smooth at the boundaries. Similarly, we impose periodic boundary condition on the second derivative of $u$.
(2) Dirichlet boundary conditions for $\mathbf{Q}$ [18, 32, 39] and natural boundary condition for $u$ are specified as follows,
(9) $\begin{cases}\mathbf{Q}=\mathbf{Q}_{bc}\text{ on }\partial\Omega,\\ \left(D^{2}u+q^{2}\left(\frac{\mathbf{Q}}{s_{+}}+\frac{I_{3}}{3}\right)u\right)\cdot\vec{\nu}=0,\left[\nabla\cdot\left(D^{2}u+q^{2}\left(\frac{Q}{s_{+}}+\frac{I_{3}}{3}\right)u\right)\right]\cdot\vec{\nu}=0,\text{ on }\partial\Omega,\end{cases}$
with the specified Dirichlet boundary $\mathbf{Q}_{bc}\in W^{\frac{1}{2},2}_{\partial\Omega,S_{0}}$, where $W^{\frac{1}{2},2}_{\partial\Omega,S_{0}}$ is a fractional order Sobolev space which is the image space of the trace operator on $W^{1,2}_{\Omega,S_{0}}$ [40]. One admissible example is the tangential Dirichlet boundary conditions in [34], for which the nematic director is tangent or in the plane of the domain boundary and such boundary conditions are motivated by experiments [17, 41]. The natural boundary condition for $u$ implies that the molecular density distribution is unconstrained on the boundary.
(3) We can also use weak boundary conditions or surface energies for the LdG order parameter as shown below [42], and the total energy is
(10) $\tilde{E}(\mathbf{Q},u)=E(\mathbf{Q},u)+\omega\int_{\partial\Omega}\|\mathbf{Q}-\mathbf{Q}_{bc}\|^{2}\mathrm{d}S,\text{ }\omega\geqslant 0,$
where $\omega\geqslant 0$ is the penalty strength. From the method of variations, the critical point of (10) satisfies the weak anchoring boundary conditions for $\mathbf{Q}$ [43, 44, 45] and natural boundary condition for $u$,
(11) $\begin{cases}\frac{\partial{\mathbf{Q}}}{\partial\vec{\nu}}+\frac{2\omega}{K}\left(\mathbf{Q}-\mathbf{Q}_{bc}\right)=0\text{ on }\partial\Omega\\ \left(D^{2}u+q^{2}\left(\frac{\mathbf{Q}}{s_{+}}+\frac{I_{3}}{3}\right)u\right)\cdot\vec{\nu}=0,\left[\nabla\cdot\left(D^{2}u+q^{2}\left(\frac{\mathbf{Q}}{s_{+}}+\frac{I_{3}}{3}\right)u\right)\right]\cdot\vec{\nu}=0,\text{ on }\partial\Omega.\end{cases}$
2.2 The proofs of existence and regularity
Proposition 2.1.
The mLdG energy functional (1) has at least a global minimiser $(\tilde{\mathbf{Q}},\tilde{u})$ in $W^{1,2}_{\Omega,\mathbf{S}_{0}}\times W^{2,2}_{\Omega}$, subject to the three types of boundary conditions enumerated above.
Proof 2.2.
The admissible space $W^{1,2}_{\Omega,\mathbf{S}_{0}}\times W^{2,2}_{\Omega}$ is non-empty. The existence of a global minimiser of (1) under Dirichlet boundary conditions for both $\mathbf{Q}$ and $u$ has been proven in [35]. We prove that the existence of a global minimiser also holds with weak anchoring for $\tilde{\mathbf{Q}}$ and natural boundary condition for $\tilde{u}$. The bulk energy $f_{bn}(\mathbf{Q})$ is a fourth-order polynomial of $\mathbf{Q}$, and the fourth-order term is positive because $C>0$. Hence, there exists a positive $M$ (that depends on $A$, $B$, $C$) such that $f_{bn}(\mathbf{Q})\geqslant\frac{C}{8}|\mathbf{Q}|^{4}$ for $|\mathbf{Q}|^{2}\geqslant M$, so that
(12) $f_{bn}(\mathbf{Q})\geqslant\begin{cases}\frac{C}{8}|\mathbf{Q}|^{4}\geqslant\frac{MC}{8}|\mathbf{Q}|^{2},|\mathbf{Q}|^{2}\geqslant M,\\ \min_{|\mathbf{Q}|^{2}\leqslant M}f_{bn}(\mathbf{Q})=constant,|\mathbf{Q}|^{2}\leqslant M.\end{cases}$
Thus, there exist two positive constants, $C_{1}(A,B,C)>0,C_{2}(A,B,C)>0$, such that
(13) $\int_{\Omega}f_{bn}(\mathbf{Q})\mathrm{d}\mathbf{x}\geqslant C_{1}(A,B,C)\|Q\|^{2}_{L^{2}_{\Omega,\mathbf{S}_{0}}}-C_{2}(A,B,C),$
and
(14) $\int_{\Omega}\frac{K}{2}\left|\nabla\mathbf{Q}\right|^{2}+f_{bn}(\mathbf{Q})\mathrm{d}\mathbf{x}+\omega\int_{\partial\Omega}\|\mathbf{Q}-\mathbf{Q}_{bc}\|^{2}\mathrm{d}S$
$\geqslant\min\left(\frac{K}{2},C_{1}(A,B,C)\right)\|Q\|_{W^{1,2}_{\Omega,\mathbf{S}_{0}}}^{2}-C_{2}(A,B,C),$
which means (1) is coercive with respect to $\mathbf{Q}$. Now we prove the coerciveness estimate in $u$, i.e. if the sequence $E(\mathbf{Q}_{i},u_{i})$ is bounded, then $u_{i}$ is also bounded in $W^{2,2}_{\Omega}$. The bulk energy $f_{bs}(u_{i})$ is a fourth order polynomial of $u_{i}$ with $c>0$, and $\int_{\Omega}f_{bs}(u_{i})\mathrm{d}\mathbf{x}$ is bounded, so $\|u_{i}\|_{L^{2}_{\Omega}}$, $\|u_{i}^{2}\|_{L^{2}_{\Omega}}$ are also bounded. Similarly, $\|\mathbf{Q}_{i}^{2}\|_{L^{2}_{\Omega,\mathbf{S}_{0}}}$ are shown to be bounded.
When $A\geqslant\frac{B^{2}}{27C}$, the boundedness of $\|D^{2}u_{i}\|^{2}_{L^{2}_{\Omega}}$ can be directly obtained from (5). For $A<\frac{B^{2}}{27C}$, we can show that $\|D^{2}u_{i}\|^{2}_{L^{2}_{\Omega}}$ is bounded by using the following inequality: $f_{1}^{2}\leqslant 2(f_{1}+f_{2})^{2}+2f_{2}^{2}$
(15) $\int_{\Omega}\left|D^{2}u_{i}\right|^{2}\mathrm{d}\mathbf{x}\leqslant\int_{\Omega}2\left|D^{2}u_{i}+q^{2}\left(\frac{\mathbf{Q}_{i}}{s_{+}}+\frac{\mathbf{I}_{3}}{3}\right)u_{i}\right|^{2}\mathrm{d}\mathbf{x}+2\int_{\Omega}\left|q^{2}\left(\frac{\mathbf{Q}_{i}}{s_{+}}+\frac{\mathbf{I}_{3}}{3}\right)u_{i}\right|^{2}\mathrm{d}\mathbf{x}.$
Given the boundedness of both $\|u_{i}\|^{2}_{L^{2}_{\Omega}}$ and $\|D^{2}u_{i}\|_{L^{2}_{\Omega}}$ along with the following inequality,
(16) $\|u_{i}\|^{2}_{L^{2}_{\Omega}}+\|D^{2}u_{i}\|_{L^{2}_{\Omega}}\geqslant C_{3}(\Omega)\|\nabla u_{i}\|^{2}_{L^{2}_{\Omega}}$
from Theorem 5.19 of [46], we have established the boundedness of $\|u_{i}\|_{W^{2,2}_{\Omega}}$ which proves the coerciveness estimate for $u$. The weak lower semi-continuity of the LdG energy and the surface energy is guaranteed in [43] and the weak lower semi-continuity features of $f_{bs}$ and N-S coupling term are guaranteed in [35]. Thus, the mLdG energy in (1) is weakly lower semi-continuous, and the existence of a global minimiser follows from the direct methods in the calculus of variations.
For $A<\frac{B^{2}}{27C}$, the Euler-Lagrange equations for the critical points of the free energy (1) are given by
(17) $K\Delta\mathbf{Q}=$ $A\cdot\mathbf{Q}-B\cdot\left(\mathbf{Q}^{2}-\frac{tr(\mathbf{Q}^{2})}{3}\mathbf{I}\right)+C\cdot tr(\mathbf{Q}^{2})\mathbf{Q}$
$+2B_{0}\cdot q^{2}/s_{+}\cdot\left(u\cdot D^{2}u-\frac{tr(u\cdot D^{2}u)}{3}\mathbf{I}\right)+2\cdot B_{0}\cdot q^{4}\cdot\frac{\mathbf{Q}}{s_{+}^{2}}u^{2},$
$2B_{0}\Delta^{2}u=$ $-au-bu^{2}-cu^{3}-4B_{0}\cdot D^{2}u:\left(q^{2}\cdot\left(\frac{\mathbf{Q}}{s_{+}}+\frac{\mathbf{I}}{3}\right)\right)$
$-2B_{0}\cdot\nabla\cdot\left(\nabla\cdot\left(q^{2}\cdot\left(\frac{\mathbf{Q}}{s_{+}}+\frac{\mathbf{I}}{3}\right)u\right)\right)-4B_{0}\cdot\left|q^{2}\cdot\left(\frac{\mathbf{Q}}{s_{+}}+\frac{\mathbf{I}}{3}\right)\right|^{2}u.$
where $\Delta^{2}u=\left(\frac{\partial^{2}}{\partial x^{2}_{1}}+\frac{\partial^{2}}{\partial x^{2}_{2}}+\frac{\partial^{2}}{\partial x^{2}_{3}}\right)^{2}u$, and we prove that the weak solutions of (17), $\bar{\mathbf{Q}}\in W_{\Omega,S_{0}}^{1,2}$,$\bar{u}\in W_{\Omega}^{2,2}$, are in fact, classical solutions of (17).
Proposition 2.3.
Let $\Omega$ be a bounded, connected open set in $\mathbb{R}^{3}$, $\partial\Omega$ is $C^{4,1/2}$ continuous, and $K,B_{0}\neq 0$, then the weak solutions $\bar{\mathbf{Q}}\in W_{\Omega,S_{0}}^{1,2},\bar{u}\in W_{\Omega}^{2,2}$ of (17) are classical solutions of (17), i.e. $\bar{\mathbf{Q}}\in C_{\Omega,S_{0}}^{2}$ and $\bar{u}\in C_{\Omega}^{4}$.
Proof 2.4.
Assume that $\bar{\mathbf{Q}}\in W_{\Omega,S_{0}}^{1,2},\bar{u}\in W_{\Omega}^{2,2}$ are weak solutions of the following Euler-Lagrange equation,
(18) $K\Delta\bar{\mathbf{Q}}=$ $\underbrace{A\cdot\bar{\mathbf{Q}}-B\cdot\left(\bar{\mathbf{Q}}^{2}-\frac{tr(\bar{\mathbf{Q}}^{2})}{3}\mathbf{I}\right)+C\cdot tr(\bar{\mathbf{Q}}^{2})\bar{\mathbf{Q}}}_{f_{1}(\bar{\mathbf{Q}})}$
$+\underbrace{2\cdot B_{0}\cdot q^{2}/s_{+}\cdot\left(\bar{u}\cdot D^{2}\bar{u}-\frac{tr(\bar{u}\cdot D^{2}\bar{u})}{3}\mathbf{I}\right)}_{f_{2}(\bar{u})}+\underbrace{2\cdot B_{0}\cdot q^{4}\cdot\frac{\bar{\mathbf{Q}}}{s_{+}^{2}}\cdot\bar{u}}_{f_{3}(\bar{\mathbf{Q}},\bar{u})},$
$\Delta^{2}\bar{u}=$ $\underbrace{-\frac{a}{2B_{0}}\bar{u}-\frac{b}{2B_{0}}\bar{u}^{2}-\frac{c}{2B_{0}}\bar{u}^{3}}_{f_{4}(\bar{u})}-\underbrace{2D^{2}\bar{u}:\left(q^{2}\cdot\left(\frac{\bar{\mathbf{Q}}}{s_{+}}+\frac{\mathbf{I}}{3}\right)\right)}_{f_{5}(\bar{\mathbf{Q}},\bar{u})}$
$-\underbrace{\nabla\cdot\left(\nabla\cdot\left(q^{2}\cdot\left(\frac{\bar{\mathbf{Q}}}{s_{+}}+\frac{\mathbf{I}}{3}\right)\bar{u}\right)\right)}_{f_{6}(\bar{\mathbf{Q}},\bar{u})}-\underbrace{2\left|q^{2}\cdot\left(\frac{\bar{\mathbf{Q}}}{s_{+}}+\frac{\mathbf{I}}{3}\right)\right|^{2}\bar{u}}_{f_{7}(\bar{\mathbf{Q}},\bar{u})}.$
From the density of $C^{\infty}_{\Omega}$ in $W^{1,2}_{\Omega}$ and $W^{2,2}_{\Omega}$ [40], we can assume that the boundary data (or trace) of $\bar{u}$ and $\bar{Q}$ coincide with functions in $C^{\infty}_{\Omega}$.
Recall that we are working in 3D case. By using the Sobolev embedding theorem in the 3D case [40], we have
(19) $u\in W_{\Omega}^{2,2}\hookrightarrow C^{0,\frac{1}{2}}_{\Omega},\mathbf{Q}\in W_{\Omega,\mathbf{S}_{0}}^{1,2}\hookrightarrow L^{6}_{\Omega,\mathbf{S}_{0}},$
and then
(20) $f_{1}(\bar{\mathbf{Q}}),f_{2}(\bar{u}),f_{3}(\bar{\mathbf{Q}},\bar{u})\in L^{2}_{\Omega,\mathbf{S}_{0}}.$
The right-hand side of the first partial differential equation is in $L^{2}_{\Omega,\mathbf{S}_{0}}$, and elliptic regularity yields $\mathbf{Q}\in W_{\Omega,\mathbf{S}_{0}}^{2,2}$, which is allowed by the regularity of boundary data and that of the domain [47]. Hence, we have
(21) $f_{4}(\bar{u})\in C^{0,\frac{1}{2}}_{\Omega}\in L^{2}_{\Omega},f_{5}(\bar{\mathbf{Q}},\bar{u}),f_{6}(\bar{\mathbf{Q}},\bar{u}),f_{7}(\bar{\mathbf{Q}},\bar{u})\in L^{2}_{\Omega}.$
Then the right-hand side of the second partial differential equation in (18) is in $L^{2}_{\Omega,\mathbf{S}_{0}}$, and elliptic regularity yields $u\in W_{\Omega}^{4,2}$. Then, the right-hand side of the first equation of (18) belongs to $W^{2,2}_{\Omega,\mathbf{S}_{0}}\hookrightarrow C^{0,1/2}_{\Omega,\mathbf{S}_{0}}$, and the Schauder estimate [48] gives $\mathbf{Q}\in C^{2,1/2}_{\Omega,\mathbf{S}_{0}}$. One can continue to alternately increase the regularity of $\bar{\mathbf{Q}}$ and $\bar{u}$ to obtain the full regularity.
Remark 2.5.
Proposition 2.3 provides the critical point with sufficient smoothness, a key feature that will be utilized in the next sections.
In the subsequent discussion, we will focus on the 2D case to facilitate comparisons with the experimental observations of smectic phases on square domains [16] and with the numerical results for nematic phases on 2D domains [29, 18]. The results in Sections 3, 4.1, 4.2, and the numerical method in Appendix can be generalized to 3D cases, by employing the same methodology.
3 Thermotropic phase transition
We consider the I-N-S phase transition with periodic boundary conditions. Consider the domain $\Omega=[0,h]$ along $x$-axis and assume that the rLdG order parameter, $\mathbf{Q}$, is of the form
(22) $\mathbf{Q}=\begin{pmatrix}Q&0\\ 0&-Q\end{pmatrix}.$
This corresponds to constant director along the $x$-axis with only one degree of freedom, the scalar order parameter $Q$ that measures the degree of ordering about the director. When $A<0$, the free energy (1) with $d=2$ is simplifies to
(23) $E_{1D}(Q,u)=\int_{0}^{h}f_{bs}(u)+B_{0}\left[u_{xx}+q^{2}\left(\frac{Q}{\sqrt{-2A/C}}+\frac{1}{2}\right)u\right]^{2}+KQ_{x}^{2}+AQ^{2}+CQ^{4}\text{ }\mathrm{d}x,$
and for $A\geqslant 0$,
(24) $E_{1D}(Q,u)=\int_{0}^{h}f_{bs}(u)+B_{0}u_{xx}^{2}+KQ_{x}^{2}+AQ^{2}+CQ^{4}\text{ }\mathrm{d}x.$
The two temperature-dependent parameters are $A=\alpha_{1}(T-T_{1}^{*})$, and $a=\alpha_{2}(T-T_{2}^{*})$, where $T_{2}^{*}<T_{1}^{*}$. It is known that the isotropic phase loses stability for $T>T_{1}^{*}$ and we show that the nematic phase (with $u=0$) loses stability at $T=T_{2}^{*}$ and the smectic phase (with non-zero $u$) is the energy minimiser for $a<0$. Hence, $T_{1}^{*}$ and $T_{2}^{*}$ are the critical temperatures for the I-N and N-S phase transitions respectively, with $T_{2}^{*}<T_{1}^{*}$ [1]. The admissible spaces are
(25) $\begin{cases}Q\in V_{Q}=\{\mathbf{Q}\in W_{\Omega}^{1,2},Q(0)=Q(h),D_{x}Q(0)=D_{x}Q(h)\},\\ u\in V_{u}=\{u\in W_{\Omega}^{2,2},u(0)=u(h),D_{x}u(0)=D_{x}u(h),D_{xx}u(0)=D_{xx}u(h)\},\end{cases}$
and the E-L equations for $A<0$ are
(26) $\begin{cases}2KQ_{xx}=2AQ+4CQ^{3}+\frac{2B_{0}q^{2}uu_{xx}}{\sqrt{-2A/C}}+\frac{2B_{0}q^{4}}{\sqrt{-2A/C}}\left(\frac{Q}{\sqrt{-2A/C}}+\frac{1}{2}\right)u^{2},\\ -2B_{0}u_{xxxx}=au+cu^{3}+4B_{0}q^{2}\left(\frac{Q}{\sqrt{-2A/C}}+\frac{1}{2}\right)u_{xx}+2B_{0}q^{2}\frac{Q_{xx}u}{\sqrt{-2A/C}}\\ \qquad\qquad\qquad+4B_{0}q^{2}\frac{Q_{x}u_{x}}{\sqrt{-2A/C}}+2B_{0}q^{4}\left(\frac{Q}{\sqrt{-2A/C}}+\frac{1}{2}\right)^{2}u.\end{cases}$
Proposition 3.1.
For any positive $c,B_{0},K,C$, and $q=\frac{2\pi n_{0}}{h},n_{0}=1,2,3,\cdots$, as temperature deceases, the energy functional (23) exhibits second-order I-N phase transition at $T=T_{1}^{*}$, and nematic phase is stable for $T_{2}^{*}\leqslant T<T_{1}^{*}$, but loses stability when $T<T_{2}^{*}$.
Proof 3.2.
The isotropic phase $(Q_{I}\equiv 0,u_{I}\equiv 0)$ is always a solution of the E-L equation of (23) for $A<0$ and (24) for $A\geqslant 0$, and the nematic phase $(Q_{N}\equiv\sqrt{-A/2C},u_{N}\equiv 0)$ is the solution of (26) when $A<0$.
For $T\geqslant T_{1}^{*}$, we have $a=\alpha_{2}(T-T_{2}^{*})\geqslant 0$, $A=\alpha_{1}(T-T_{1}^{*})\geqslant 0$ i.e. $f_{bs}(u)\geqslant 0,AQ^{2}+CQ^{4}\geqslant 0$. Hence, for any $Q,u$ in admissible space, the isotropic phase $(Q_{I}\equiv 0,u_{I}\equiv 0)$ is the global minimiser for $T\geqslant T_{1}^{*}$, i.e.
(27) $E_{1D}(Q,u)=\int_{0}^{h}f_{bs}(u)+B_{0}u_{xx}^{2}+KQ_{x}^{2}+AQ^{2}+CQ^{4}\text{ }\mathrm{d}x\geqslant 0=E_{1D}(Q_{I},u_{I}).$
For $T_{1}^{*}>T\geqslant T_{2}^{*}$, we have $a=\alpha_{2}(T-T_{2}^{*})\geqslant 0$, i.e. $f_{bs}(u)\geqslant 0$. Hence, for any $Q,u$ in admissible space, nematic phase $(Q_{N}\equiv\sqrt{-A/2C},u_{N}\equiv 0)$ is the global minimiser for $T_{1}^{*}>T\geqslant T_{2}^{*}$, since
(28) $E_{1D}(Q,u)=\int_{0}^{h}$ $f_{bs}(u)+B_{0}\left[u_{xx}+q^{2}\left(\frac{Q}{\sqrt{-2A/C}}+\frac{1}{2}\right)u\right]^{2}$
$+KQ_{x}^{2}+AQ^{2}+CQ^{4}\text{ }\mathrm{d}x$
$\geqslant\int_{0}^{h}$ $AQ^{2}+CQ^{4}\text{ }\mathrm{d}x\geqslant\int_{0}^{h}-\frac{A^{2}}{4C}\text{ }\mathrm{d}x=E_{1D}(Q_{N},u_{N}).$
To investigate the stability of nematic phase near $T=T^{*}_{2}$, we calculate the second variation of (23) at $(Q_{N}\equiv\sqrt{-A/2C},u_{N}\equiv 0)$ along period perturbation $(\eta_{1},\eta_{2})$,
(29) $\delta^{2}E_{1D}(\eta_{1},\eta_{2})=\int_{0}^{h}\left(a(T)\cdot\eta_{2}^{2}+2B_{0}\left(\eta_{2xx}+q^{2}\eta_{2}\right)^{2}+2K(\eta_{1x})^{2}-4A\eta_{1}^{2}\right)\mathrm{d}x.$
The stability of the nematic phase is measured by the minimum eigenvalue of $\delta^{2}E_{1D}$, i.e.
(30) $\mu_{T}=\inf_{\eta_{1}\in V_{Q},\eta_{2}\in V_{u}}\frac{\delta^{2}E_{1D}(\eta_{1},\eta_{2})}{\int_{0}^{h}\eta_{1}^{2}+\eta_{2}^{2}\mathrm{d}x}.$
If $\mu_{T}<0$, the nematic phase is unstable. If $\mu_{T}>0$, the nematic phase is stable.
For $T<T^{*}_{1}$, i.e., $-4A>0$, any perturbation with a non-zero $\eta_{1}$ is always a stable direction. Thus, we only consider the perturbation $(0,\eta_{2})$. The Fourier expansion of $\eta_{2}$ in $\Omega=[0,h]$ is given by
(31) $\eta_{2}=w_{0}/2+\sum_{n=1}^{\infty}w_{n}\cos(\frac{2\pi nx}{h})+v_{n}\sin(\frac{2\pi nx}{h}).$
By substituting (31) into (29), we have
(32) $\delta^{2}E_{1D}(0,\eta_{2})=h/2\cdot\left(\frac{a+2B_{0}q^{4}}{2}\right)w_{0}^{2}+h/2\cdot\sum_{n=1}^{\infty}\left[2B_{0}\left(\frac{4\pi^{2}n^{2}}{h^{2}}-q^{2}\right)^{2}+a\right](w_{n}^{2}+v_{n}^{2}).$
$(0,\eta)$ is an eigenfunction of $E_{1D}$ at nematic phase if and only if
(33) $2a\eta+4B_{0}\eta_{xxxx}+8B_{0}q^{2}\eta_{xx}+4B_{0}q^{4}\eta=\lambda\eta.$
One can verify that (33) is the first order optimal condition (or KKT condition [49]) of (30). By substituting (31) into (33), we get that $\eta\equiv 1$, $\eta=\cos(\frac{2\pi nx}{h})$ and $\eta=\sin(\frac{2\pi nx}{h})$, $n=1,2,3\cdots$ are the eigenvectors of $\delta^{2}E$ with eigenvalues $\mu=a+2B_{0}q^{4}$ and $a+2B_{0}\left(\frac{4\pi^{2}n^{2}}{h^{2}}-q^{2}\right)^{2}$, $n=1,2,3\cdots$, respectively. If there is a $n\in\mathbb{Z^{+}}$ s.t. $\left(\frac{4\pi^{2}n^{2}}{h^{2}}-q^{2}\right)^{2}=0$, then $\eta=\sin(\frac{2\pi nx}{h})=\sin\left(qx\right)$ and $\eta=\cos(\frac{2\pi nx}{h})=\cos\left(qx\right)$ are the eigenvectors corresponding to the minimum degenerate eigenvalue $\mu=a$. For $T_{1}^{*}>T\geqslant T_{2}^{*}$, i.e., $a\geqslant 0$, the second variation is always positive, i.e., the nematic phase is stable. For $T<T_{2}^{*}$, i.e., $a<0$, the eigenvector $\eta\equiv 1$ is an unstable eigendirection if and only if the corresponding eigenvalue $a+2B_{0}q^{4}<0$ is negative, and the eigenvectors $\sin(\frac{2\pi nx}{h})$ and $\cos(\frac{2\pi nx}{h})$, $n=1,2,3\cdots$, are unstable eigendirections if and only if the corresponding eigenvalue $a+2B_{0}\left(\frac{4\pi^{2}n^{2}}{h^{2}}-q^{2}\right)^{2}$ is negative. Thus, the Morse index of the nematic phase, i.e., the number of eigenvectors corresponding to negative eigenvalues is
(34) $i_{nematics}=2\times card(\mathbb{N}_{nematics})+n_{0},$
where
(35) $\mathbb{N}_{nematics}=\left\{n\in\mathbb{Z^{+}}:a+2B_{0}\left(\frac{4\pi^{2}n^{2}}{h^{2}}-q^{2}\right)^{2}<0\right\}$
and $card(\mathbb{N}_{nematics})$ is the cardinal number of $\mathbb{N}_{nematics}$. If $a+2B_{0}q^{4}\geqslant 0$, $n_{0}=0$; otherwise $n_{0}=1$, i.e. $\eta\equiv 1$ is an unstable eigendiretion. As $a<0$ decreases, more positive integers satisfy the constraint in (35), and the Morse index of the nematic phase $i_{nematics}$ increases.
For example, in Figure 1, we substitute the parameter values in the caption to (34), and get =$\mathbb{N}=\{3,4,5\}$ and $n_{0}=0$, i.e., the Morse index of the nematic phase $i_{nematics}=6$ with unstable eigendirections $\eta=\sin(nx),\cos(nx)$, $n\in\mathbb{N}$.
Figure 1: The nematic critical point for $h=2\pi$, $q=4$, $T=-30$, $T_{1}^{*}=0$, $T_{2}^{*}=-10$, $a=T-T_{2}^{*}=-20$, $A=T-T_{1}^{*}=-30$, $B_{0}=0.1$, $c=10,C=10$. $V_{1}$ to $V_{6}$ are the unstable eigendirections associated with $u$ and $V_{7}$ is a stable eigendirection. The pairs of unstable eigendirections $V_{1}$ and $V_{2}$, $V_{3}$ and $V_{4}$, $V_{5}$ and $V_{6}$ are the orthogonal linear combinations of $\sin(nx)$ and $\cos(nx)$ with $n=4,3,5$ respectively. The colour bar represents the modulation of the concentration, the white lines define the nematic director: (±$(\pm 1,0)$ in all subsequent figures.
In the above, we prove that the nematic phase loses stability as temperature decreases. In the remaining part, we demonstrate that when the nematic phase loses stability, it bifurcates into a more stable smectic phase. To study this, we consider the following E-L equation for $u$,
(36) $2B_{0}u_{xxxx}+au+cu^{3}+4B_{0}q^{2}u_{xx}+2B_{0}q^{4}u=0,$
i.e. fix $Q\equiv\frac{s_{+}}{2}$ in (26) for brevity, but the results also hold for variable $Q$. In the proof of Proposition 3.1, we note that the minimum eigenvalue of the nematic phase is degenerate, which presents technical difficulties in bifurcation theory [50]. To circumvent this issue, we construct the following working space:
(37) $V=V_{u}\cap W^{1,2}_{0,\Omega},$
where $V_{u}$ is defined in (25). This restricts $\eta=cos(qx)$ from serving as an eigenvector and then simplifies the minimum eigenvalue at the nematic phase.
Proposition 3.3.
For any positive $c,B_{0}$, and $q=\frac{2\pi n_{0}}{h},n_{0}=1,2,3,\cdots$, a pitchfork bifurcation of (36) arises at $a=0$ or $T=T_{2}^{*}$, $u\equiv 0$ in $V$. More precisely, there exists positive numbers $\epsilon,\delta$ and two smooth maps
(38) $t\in(-\delta,\delta)\rightarrow a(t)\in(-\epsilon,\epsilon),t\in(-\delta,\delta)\rightarrow w_{t}\in V$
such that all the pairs $(a,u)\in R\times V$ satisfying
[…]
are either
 $nematic\text{ }phase:u\equiv 0\text{ }or\text{ }smectic\text{ }phases:u=\pm\left(tsin(qx)+t^{2}w_{t}\right).$
Proof 3.4.
The proof follows the same paradigm as in Theorem 5.2 in [51] and Theorem 5.1 in [29], and we address the necessary technical differences that arise because the study in [51] and [29] focuses on a second-order partial differential equation, while our analysis involves a fourth-order partial differential equation.
To show that a pitchfork bifurcation arises at $a=0$, we apply the Crandall and Rabinowitz bifurcation theorem [52] to the operator :×$\mathcal{F}:\mathbb{R}\times V\rightarrow W^{-2,2}_{\Omega}$ ($W^{-2,2}_{\Omega}$ is the dual space of $W^{2,2}_{\Omega}$) defined by
(39) $\mathcal{F}(a,w):=2B_{0}D_{xxxx}w+aw+cw^{3}+4B_{0}q^{2}D_{xx}w+2B_{0}q^{4}w.$
We have to check four assumptions of Theorem 1.7 in [52]:
(a) $\mathcal{F}(a,0)=0$; (b) The partial derivatives $D_{a}\mathcal{F},D_{w}\mathcal{F},D_{aw}\mathcal{F}$ exist and are continuous; (c)$dim\left(\frac{W^{-2,2}(\Omega)}{Range(D_{w}\mathcal{F}(0,0))}\right)=dim\left(Kernel\left(D_{w}\mathcal{F}(0,0)\right)\right)=1$; (d) $D_{aw}\mathcal{F}w_{0}\notin Range(D_{w}\mathcal{F}(0,0))$, where $w_{0}\in Kernel\left(D_{w}\mathcal{F}(0,0)\right)$.
$\mathcal{F}(a,0)=0$ holds for all $a\in\mathbb{R}$. We have
(40) $\begin{cases}D_{a}\mathcal{F}(a,w)=w,\\ D_{w}\mathcal{F}(a,w)=2B_{0}D_{xxxx}+a+3cw^{2}+4B_{0}q^{2}D_{xx}+2B_{0}q^{4},\\ D_{aw}\mathcal{F}(a,w)=1,\end{cases}$
and they are continuous, since $D_{w}\mathcal{F}(a,w):V\rightarrow W^{-2,2}_{\Omega}$ is a bounded linear operator. For checking $\mathcal{F}$ satisfies assumption (c), we should calculate the kernel space of
(41) $D_{w}\mathcal{F}(0,0)=2B_{0}D_{xxxx}+4B_{0}q^{2}D_{xx}+2B_{0}q^{4}=2B_{0}(D_{xx}+q^{2})(D_{xx}+q^{2})$
in $V$, i.e. the solution space of the following differential equation:
(42) $\begin{cases}D_{w}\mathcal{F}(0,0)w=2B_{0}D_{xxxx}w+4B_{0}q^{2}D_{xx}w+2B_{0}q^{4}w=0,\\ w(0)=w(h)=0,D_{x}w(0)=D_{x}w(h),D_{xx}w(0)=D_{xx}w(h).\end{cases}$
The general solution of the differential question in (42) without considering the boundary condition is
(43) $w=(k_{1}+k_{2}x)sin(qx)+(k_{3}+k_{4}x)cos(qx),k_{i}\in\mathbb{R},i=1,2,3,4.$
By taking the boundary condition into account, we have $w=k_{1}sin(qx),k_{1}\in\mathbb{R}$, and
(44) $dim\left(Kernel\left(D_{w}\mathcal{F}(0,0)\right)\right)=dim\left(\{w=k_{1}sin(qx),k_{1}\in\mathbb{R}\}\right)=1.$
For any $u_{a},u_{b}\in V$, we have
(45) $D_{w}\mathcal{F}(0,0)u_{a},u_{b}\right>$ $=2B_{0}\int_{0}^{h}\left((D_{xx}+q^{2})(D_{xx}+q^{2})u_{a}\right)u_{b}\mathrm{d}x$
$=2B_{0}\int_{0}^{1}\left((D_{xx}+q^{2})(D_{xx}+q^{2})u_{b}\right)u_{a}\mathrm{d}x=\left<u_{a},D_{w}\mathcal{F}(0,0)u_{b}\right>$
by using the boundary conditions of $u_{a}$ and $u_{b}$, which means $D_{w}\mathcal{F}(0,0)$ is a self-adjoint operator, and hence it is a Fredholm operator of index 00 [53], and we have
(46) $dim\left(\frac{W^{-2,2}(\Omega)}{Range(D_{w}\mathcal{F}(0,0))}\right)=dim\left(Kernel\left(D_{w}\mathcal{F}(0,0)\right)\right)=1,$
which satisfies assumption (c). We also need to check the last assumption (d),
(47) $D_{aw}\mathcal{F}(a,w)sin(qx)=sin(qx)\notin rangeD_{w}\mathcal{F}(0,0),$
i.e. the following differential equation,
(48) $\begin{cases}2B_{0}D_{xxxx}w+4B_{0}q^{2}D_{xx}w+2B_{0}q^{4}w=sin(qx),\\ w(0)=w(h)=0,D_{x}w(0)=D_{x}w(h)=0,D_{xx}w(0)=D_{xx}w(h),\end{cases}$
does not have a solution. One can check that the general solution of (48) without considering the boundary condition is
(49) $w=-\frac{x^{2}sin(qx)}{16B_{0}q^{2}}+k_{1}sin(qx),k_{1}\in\mathbb{R},$
and it can not satisfy the boundary conditions with any $k_{1}\in\mathbb{R}$, so that $sin(qx)\notin Range\left(D_{w}\mathcal{F}(0,0)\right)$. All the assumptions of Crandall and Rabinowitz’s theorem are satisfied, and our proposition now follows directly by Crandall and Rabinowitz’s result [52].
Remark 3.5.
In Proposition 4, we fix $Q\equiv\frac{s_{+}}{2}$ in (26) for brevity, but the results also hold for coupled system (26) without treating $Q$ to be a constant by defining $\mathcal{F}(a,w_{1},w_{2})=(\mathcal{F}_{1}(a,w_{1},w_{2}),\mathcal{F}_{2}(a,w_{1},w_{2})):\mathbb{R}\times V_{Q}\times V\rightarrow W^{-1,2}_{\Omega}\times W^{-2,2}_{\Omega}$ where
[…]
$A(a)=\alpha_{1}(\frac{a}{\alpha_{2}}+T_{2}^{*}-T_{1}^{*})$ and $s_{+}(a)=\sqrt{-2A(a)/C}$. One can check that
(51) $D_{(w_{1},w_{2})}\mathcal{F}(0,0,0)=\left(-2KD_{xx}-4A(0),2B_{0}(D_{xx}+q^{2})(D_{xx}+q^{2})\right)$
is also a Fredholm operator of index 0, and $dim(Kernel(D_{(w_{1},w_{2})}\mathcal{F}(0,0,0))=1$ since the spectrum [53] of $-2KD_{xx}-4A(0),A(0)<0$ in $V_{Q}$ is positive which does not change the dimension of kernel space. Thus, we can directly have the following proposition.
Proposition 3.6.
For any positive $c,B_{0},K,C$, and $q=\frac{2\pi n_{0}}{h},n_{0}=1,2,3,\cdots$ in (26), the nematic phase ($Q\equiv s_{+}/2,u\equiv 0$) loses stability in $V_{Q}\times V$ at the critical temperature $T=T_{2}^{*}$ (or $a=0$), and this loss of stability is accompanied by a symmetric pitchfork bifurcation.
In Figure 2, we numerically calculate (the numerical method is specified in Appendix) the N-S bifurcation, accomplished using the sine spectral method for $u$ [54] and Fourier spectral method for $Q$ [55]. This numerical scheme covers the boundary conditions in $V_{Q}\times V$. For $a>0$, the minimum eigenvalue at the nematic phase, as calculated both numerically and analytically, is both simple and positive, indicating stability. When $a=0$, a simple zero eigenvalue emerges with eigenvector $\eta=\sin(qx)$. As $a$ becomes negative, the nematic phase loses stability and bifurcates into two smectic phases, corresponding to $u=t\sin(qx)+t^{2}w_{t}$ and $u=-t\sin(qx)-t^{2}w_{t}$ respectively, in pitchfork bifurcation.
The numerically calculated bifurcation diagram of the I-N-S phase transition v.s. temperature $T$ is shown in Figure 3. The isotropic phase with $u_{I}\equiv 0$ and $\mathbf{Q}_{I}\equiv\mathbf{0}$ is always a critical solution of (23). When $T\geqslant T_{1}^{*}$, the isotropic phase is a global minimiser of (23). For $T_{1}^{*}>T\geqslant T_{2}^{*}$, the isotropic phase loses stability, and the nematic phase with $u_{N}\equiv 0$ and $\mathbf{Q}_{N}\not\equiv\mathbf{0}$ becomes stable. For $T<T_{2}^{*}$, the nematic phase loses stability and the smectic phase with $u_{S}\not\equiv 0$ and $\mathbf{Q}_{S}\not\equiv\mathbf{0}$ becomes stable.
Figure 2: Schematic illustration of the the N-S phase transition with $a=T+10$, $b=0$, $c=10$, $A=T$, $C=10$, $K=0.2$ , $q=4$, $B_{0}=0.001$, and the pitchfork bifurcation for $a<0$. The solid black line denotes a stable phase, while the dashed black line denotes an unstable phase in all figures. We numerically calculate the minimiser ($u_{min}$,$Q_{min}$) of (23) with various $a$, and plot $u_{min}$. We track the bifurcation across $-5\geqslant T\geqslant-15$ ($55\geqslant a\geqslant-$ 5).
Figure 3: Phase transitions for $T_{1}^{*}=0$, $T_{2}^{*}=-10$, $\alpha_{1}=\alpha_{2}=1$, $C=c=10$, $h=2\pi,q=4$. We use $u_{max}$ and $Q_{max}$, where $u_{max}$ = $\max_{0\leqslant x\leqslant h}u(x)$ and $Q_{max}$ = $\max_{0\leqslant x\leqslant h}Q(x)$. For better visualisation, we plot the 2D $y$-invariants: $\bar{Q}(x,y)\equiv Q(x)$ and $\bar{u}(x,y)\equiv u(x)$.
4 Smectics under confinement
In this section, we focus on the low temperature regime (i.e., $a<0$ and $A<0$) to investigate smectic profiles under confinement. In Sections 4.1 and 4.2, we study the minimisers of the coupling energy, assuming a given rLdG $\mathbf{Q}$-profile, compatible with a defect-free perfectly ordered nematic state and a nematic defect respectively. These formal calculations give us some heuristic insight into how smectic layers respond to nematic profiles, with and without defects i.e. do defects repel smectic layers and do smectic layers concentrate near well-ordered nematic regions and if so, is there a correlation between the layer normal and the nematic director?
4.1 The positional order far from defects
Based on previous work [21, 56], we assume that far away from defects in confined geometries.
(52) $\mathbf{Q}=s_{+}\left(\mathbf{n}\otimes\mathbf{n}-\frac{\mathbf{I}}{2}\right)$
models a perfectly ordered nematic state, which is also a minimiser of $f_{bn}$ in (1), with arbitrary 2D nematic director $\mathbf{n}$. Based on the analysis in Section 3, we assume a simple periodic structure for $u$, compatible with a layer structure,,
(53) $u(\mathbf{x})=k_{1}cos(\tilde{q}\mathbf{k}\cdot\mathbf{x}),$
where $\mathbf{k}=\frac{\nabla u}{|\nabla u|}$, if $|\nabla u|\neq 0$, is the layer normal, and $\tilde{q}$ is the wave number of the layer. Substituting (52) and (53) into the coupling term (5), we obtain
(54) $\left|D^{2}u+q^{2}\left(\frac{\mathbf{Q}}{s_{+}}+\frac{\mathbf{I}_{2}}{2}\right)u\right|^{2}=k_{1}^{2}\left|-\tilde{q}^{2}\mathbf{k}\otimes\mathbf{k}\cdot u+q^{2}\mathbf{n}\otimes\mathbf{n}\cdot u\right|^{2}.$
The above coupling term is minimised by $\tilde{q}=q$, $\mathbf{k}=\mathbf{n}$. Thus, we deduce that away from defects, we can interpret the phenomenological parameter $q$ in (1) to be the wave number of the smectic layers and the smectic layer normal is aligned with the nematic director, $\mathbf{n}$, in perfect agreement with the definition of SmA. Of course, these deductions do not shed light into the structure of arbitrary critical points of (1).
4.2 The positional order near defects
We can assume $\mathbf{Q}\equiv 0$ near defects in the rLdG model [32]. Substituting $\mathbf{Q}=0$ into the coupling term (5), we obtain
(55) $E_{couple}(\mathbf{Q}\equiv 0,u)=\int_{\Omega}B_{0}\left|D^{2}u+\frac{q^{2}u}{2}\mathbf{I}_{2}\right|^{2}\mathrm{d}\mathbf{x}.$
It’s straightforward to verify that $u\equiv 0$ is a global minimiser since $E_{couple}(\mathbf{Q}\equiv 0,u)\geqslant 0=E_{couple}(\mathbf{Q}\equiv 0,u\equiv 0)$. Our aim is to demonstrate that $u\equiv 0$ is indeed the unique minimiser, which implies that domains with defects do not support layered structures. We prove (a) $E_{couple}(\mathbf{Q}\equiv 0,u)$ is convex, so that every minimiser $u^{*}$ is a global minimiser, i.e. $E_{couple}(u^{*})=0$, and (b) if $E_{couple}(u^{*})=0$, then $u^{*}\equiv 0$. (a) is obvious, since (55) is the $L^{2}_{\Omega}$ norm of linear $D^{2}u+\frac{q^{2}u}{2}\mathbf{I}_{2}$. Next, we prove (b). If $\left|D^{2}u+\frac{q^{2}u}{2}\mathbf{I}_{2}\right|^{2}\equiv 0$, then $u_{xy}\equiv 0$, $u_{xx}\equiv u_{yy}=-\frac{q^{2}u}{2}$. From the regularity result in Proposition 2, we can assume that $u$ has $C^{3}$ regularity. Since $u_{xy}\equiv 0$, then $u_{xxy}=-\frac{q^{2}u_{y}}{2}\equiv 0$, $u_{xyy}=-\frac{q^{2}u_{x}}{2}\equiv 0$, which imply $u_{x}=u_{y}\equiv 0$, and further $u\equiv C_{0}$ where $C_{0}$ is a constant. Then we deduce $C_{0}=0$ from $u_{xx}=u_{yy}=-\frac{q^{2}u}{2}\equiv 0$. Hence, (a) and (b) hold, which means that $u\equiv 0$ is the unique minimiser of (55).
In Figure 4, given a $\mathbf{Q}$-field on a square domain with edge length $\lambda$ and natural boundary conditions for $\mathbf{Q}$ and $u$, we plot the numerical minimiser $u$ of (1) with relatively large $B_{0}$ and relatively small $a$ and $c$. The $u$ almost vanishes at the central point defect and produces a layered structure far away from the defect, in agreement with our analysis above.
Figure 4: The coupling-energy minimising profile for $u$ for a fixed $\mathbf{Q}$-field on the left. This $\mathbf{Q}$-field has a $+1$ central point defect. The parameters are set to be: $a=-0.1$, $c=0.1$, $\lambda^{2}=30$, $q=2\pi$, $B_{0}=10^{-3}$. The colour bar of left plot is the order parameter $\sqrt{\text{Tr}(\mathbf{Q}^{2})/2}$ and the white lines model the eigenvector corresponding to the largest eigenvalue of $\mathbf{Q}$. We use the same color bar as before.
4.3 Structural transitions for smectics on square domains.
We consider qualitative properties of energy minimisers of (1) on 2D square domains: $\Omega=[-\lambda,\lambda]^{2}$. By rescaling the system according to $\bar{\mathbf{x}}=\frac{\mathbf{x}}{\lambda}$, $\bar{E}=\frac{E}{K}$, $\bar{\lambda}^{2}=\frac{2C\lambda^{2}}{K}$, $\bar{a}=\frac{a}{2C}$, $\bar{c}=\frac{c}{2C}$, $\bar{q}^{2}=\frac{Kq^{2}}{2C}$, $\bar{B}_{0}=\frac{2B_{0}C}{K^{2}}$, $\bar{A}=\frac{A}{2C}$ where the unit of $B_{0}$ is Nm2, the unit of $K$ is N, the unit of $\lambda$ is m, and the unit of $q$ is m-1. Then the non-dimensionalised energy is given by
(56) $\bar{E}(\mathbf{Q},u)=\int_{[-1,1]^{2}}$ $\left(\bar{\lambda}^{2}\left(\frac{\bar{a}}{2}u^{2}+\frac{\bar{c}}{4}u^{4}\right)+\frac{\bar{B}_{0}}{\bar{\lambda}^{2}}\left|D^{2}u+\bar{\lambda}^{2}\bar{q}^{2}\left(\frac{\mathbf{Q}}{s_{+}}+\frac{\mathbf{I}_{2}}{2}\right)u\right|^{2}\right.$
$\quad+\left.\frac{1}{2}\left|\nabla\mathbf{Q}\right|^{2}+\bar{\lambda}^{2}\left(\frac{\bar{A}}{2}\text{tr}\mathbf{Q}^{2}+\frac{(\text{tr}\mathbf{Q}^{2})^{2}}{8}\right)\right)\mathrm{d}\bar{x}.$
In the following, we drop all the bars, and the E-L equations of (56) are
(57) $\Delta\mathbf{Q}=$ $\lambda^{2}\left(A\cdot\mathbf{Q}+\frac{\text{tr}(\mathbf{Q}^{2})\mathbf{Q}}{2}\right)$
$+2B_{0}\cdot q^{2}/s_{+}\cdot\left(u\cdot D^{2}u-\frac{tr(u\cdot D^{2}u)}{2}\mathbf{I}\right)+2\lambda^{2}B_{0}\cdot q^{4}\cdot\frac{\mathbf{Q}}{s_{+}^{2}}u^{2},$
$\Delta^{2}u=$ $-\lambda^{4}\left(\frac{a}{2B_{0}}u+\frac{c}{2B_{0}}u^{3}\right)-\lambda^{2}D^{2}u:\left(q^{2}\cdot\left(\frac{\mathbf{Q}}{s_{+}}+\frac{\mathbf{I}}{2}\right)\right)$
$-\lambda^{2}\nabla\cdot\left(\nabla\cdot\left(q^{2}\cdot\left(\frac{\mathbf{Q}}{s_{+}}+\frac{\mathbf{I}}{2}\right)u\right)\right)-2\lambda^{4}\left|q^{2}\cdot\left(\frac{\mathbf{Q}}{s_{+}}+\frac{\mathbf{I}}{2}\right)\right|^{2}u.$
Regarding the boundary conditions, we assume Dirichlet tangent boundary conditions for the nematic director i.e. the director, $\mathbf{n}=\pm(1,0)$ on the horizontal edges and $\mathbf{n}=\pm(0,1)$ on the vertical edges, and the density is naturally distributed, i.e.,
(58) $\begin{cases}\mathbf{Q}=\begin{pmatrix}s_{+}L(x)/2&0\\ 0&-s_{+}L(x)/2\end{pmatrix}\text{ on }y=\{1,-1\},\\ \mathbf{Q}=\begin{pmatrix}-s_{+}L(y)/2&0\\ 0&s_{+}L(y)/2\end{pmatrix}\text{ on }x=\{1,-1\},\\ \left(D^{2}u+\lambda^{2}q^{2}\left(\frac{\mathbf{Q}}{s_{+}}+\frac{I_{2}}{2}\right)u\right)\cdot\vec{\nu}=0,\left[\nabla\cdot\left(D^{2}u+\lambda^{2}q^{2}\left(\frac{Q}{s_{+}}+\frac{I_{2}}{2}\right)u\right)\right]\cdot\vec{\nu}=0,\text{ on }\partial\Omega,\end{cases}$
where
(59) $L(x)=\begin{cases}\frac{x+1}{\epsilon_{0}},-1\leqslant x\leqslant-1+\epsilon_{0},\\ 1,|x|\leqslant 1-\epsilon_{0},\\ \frac{1-x}{\epsilon_{0}},1-\epsilon_{0}\leqslant x\leqslant 1,\end{cases}$
is a trapezoidal function with a small enough $\epsilon_{0}$, to avoid the mismatch in the boundary conditions, at the square vertices [33, 29, 57].
4.3.1 Large domain size limit
In the $\lambda\rightarrow\infty$ limit or in the Oseen-Frank limit, we can assume that the interior profile is almost a minimiser of $f_{bn}$ in (1) with no defects [34]. Analogous to the discussion in Section 4.1, we assume
(60) $\mathbf{Q}\equiv s_{+}\left(\mathbf{n}_{0}\times\mathbf{n}_{0}-\frac{\mathbf{I}}{2}\right),$
where $\mathbf{n}_{0}=(\cos\theta,\sin\theta)$ with a fixed $\theta$. We further assume a periodic structure for
(61) $u=A_{0}\cos(\mathbf{k}\cdot\mathbf{x})$
with unknown $A_{0}$ and $\mathbf{k}$, where $A_{0}$ is the amplitude of the layers, $\frac{|\mathbf{k}|}{2\pi}$ (if $|k|\neq 0$) is the wave number of layers, and $\frac{\mathbf{k}}{|\mathbf{k}|}$ is the layer normal.
By substituting (61) and (60), we have that the leading order terms in (56), in the $\lambda\to\infty$ limit are:
(62) $\lambda^{2}\int_{\Omega}\left(\frac{au^{2}}{2}+\frac{cu^{4}}{4}+\frac{B_{0}}{\lambda^{4}}\left|D^{2}u+\lambda^{2}q^{2}\left(\frac{\mathbf{Q}}{s_{+}}+\frac{\mathbf{I_{2}}}{2}\right)u\right|^{2}+\frac{A}{2}\text{tr}\mathbf{Q}^{2}+\frac{(\text{tr}\mathbf{Q}^{2})^{2}}{8}\right)\mathrm{d}\mathbf{x}$
$=\lambda^{2}\left(aA_{0}^{2}+\frac{3cA_{0}^{4}}{8}+2B_{0}A_{0}^{2}\left|q^{2}\mathbf{n}_{0}\times\mathbf{n}_{0}-\frac{\mathbf{k}\times\mathbf{k}}{\lambda^{2}}\right|^{2}+Constant+O\left(\frac{1}{|\mathbf{k}|}\right)\right).$
The leading order energy in (62) is minimised by
(63) $\mathbf{k}=q\lambda\mathbf{n}_{0},A_{0}=\sqrt{\frac{-4a}{3c}},$
since the constant can be set to zero by adding a suitable constant to $f_{bn}$ in (62). These relations contain useful information: (i) the layer normal is aligned with $\mathbf{n}_{0}$; (ii) the number of layers is proportional to $\lambda$ and the (iii) layer thickness, $l$ is inversely proportional to $q$, in the $\lambda\to\infty$ limit. Further, the amplitude of the layer oscillations, $A_{0}$, depends on the parameters of $f_{bs}$ as expected, at least for energy minimisers in the $\lambda\to\infty$ limit. In the left two plots of Figure 5, we fix $\mathbf{n}_{0}=(\sqrt{2}/2,\sqrt{2}/2)$ in (60), and numerically calculate the minimiser of $u$ in (62) without assuming the profile of $u$ in (61). In the numerical results, the wave number is proportional to $\lambda$; the layer normal follows the director $\mathbf{n_{0}}$; the amplitude of $u$ is close to the $A_{0}$ in (63). More specifically, the number of layers for $\lambda^{2}=50$ is $20$, which is equal to the predicted value $\frac{|\mathbf{k}|*4\sqrt{2}}{2\pi}=20$ (where $\frac{|\mathbf{k}|}{2\pi}$ denotes the number of layers in a unit length, and $2\sqrt{2}$ represents the length of the diagonal of the square) in (63), and the amplitude is 1.1432, close to the predicted value $A_{0}=\sqrt{\frac{-4a}{3c}}\approx 1.1547$ in (63). The number of layers for $\lambda^{2}=150$ is 35 and the predicted value is $\frac{|\mathbf{k}|*2\sqrt{2}}{2\pi}=34.6410$ in (63). The numerically calculated amplitude is 1.1403 whilst the predicted value is $A_{0}=\sqrt{\frac{-4a}{3c}}\approx 1.1547$ in (63). In the right plot of Figure 5, the director field is compatible with the boundary condition (58). The number of layers along the diagonal is also 35, and the numerically calculated amplitude is 1.1474, both of which are also closed to the predicted value.
Figure 5: The distribution of layers, u, is calculated by minimising (56) with fixed $\mathbf{Q}$-field. In the left two plots, $\mathbf{Q}_{11}=0,\mathbf{Q}_{12}=\frac{s_{+}}{2}$, i.e. the nematic director is uniformly aligned along the line $y=x$. This is not compatible with the boundary conditions in (58). In the right plot, $Q_{11}=s_{+}\cos(2\theta)/2$, $Q_{12}=s_{+}\sin(2\theta)/2$, $\theta$ is a solution of the Laplace equation, compatible with the boundary conditions in (58). The parameters are $a=-5$, $c=5$, $B_{0}=10^{-3}$, $q=2\pi$.
4.3.2 Small domain size limit
In the study of nematic configuration confined on a square, Well Order Reconstruction Solution (WORS) [58, 29, 18] with two crossed line defects is the unique stable state with small enough $\lambda$. In this subsection, we will show that in $\lambda\rightarrow 0$ limit or a very small domain size, the stable smectic state has the same $\mathbf{Q}$ profile as WORS, i.e. $\mathbf{Q}\rightarrow\mathbf{Q}_{WORS}$ and $u$ does not have a layer structure.
In the $\lambda\rightarrow 0$ limit, we can take a regular perturbation expansion of $\mathbf{Q}$ and $u$ in powers of $\lambda$ as shown below:
(64) $\mathbf{Q}=\mathbf{Q}_{0}+\lambda\mathbf{Q}_{1}+\lambda^{2}\mathbf{Q}_{2}+\cdots,u=u_{0}+\lambda u_{1}+\lambda^{2}u_{2}+\cdots$
where $(\mathbf{Q}_{0},u_{0})$ is the solution of the following partial differential equation:
(65) $\begin{cases}\Delta\mathbf{Q}_{0}=2B_{0}\cdot q^{2}/s_{+}\cdot\left(u_{0}\cdot D^{2}u_{0}-\frac{tr(u_{0}\cdot D^{2}u_{0})}{2}\mathbf{I_{2}}\right)\\ \Delta^{2}u_{0}=0\end{cases},$
which satisfies the boundary condition:
(66) $\begin{cases}\mathbf{Q}_{0}=\begin{pmatrix}s_{+}L(x)/2&0\\ 0&-s_{+}L(x)/2\end{pmatrix}\text{ on }y=\{1,-1\},\\ \mathbf{Q}_{0}=\begin{pmatrix}-s_{+}L(y)/2&0\\ 0&s_{+}L(y)/2\end{pmatrix}\text{ on }x=\{1,-1\}\\ D^{2}u_{0}\cdot\vec{\nu}=0,\left[\nabla\cdot D^{2}u_{0}\right]\cdot\vec{\nu}=0,\text{ on }\partial\Omega.\end{cases}$
Proposition 4.1.
The solutions of (65) with boundary conditions (66) are
(67) $\begin{cases}\mathbf{Q}_{0}(x,y)=\begin{pmatrix}Q_{0}(x,y)&0\\ 0&-Q_{0}(x,y)\end{pmatrix},\\ u_{0}(x,y)=k_{1}x+k_{2}y+k_{3},k_{i}\in\mathbb{R},i=1,2,3,\end{cases}$
where
(68) $Q_{0}(x,y)=$ $\sum_{k\text{ odd}}\frac{8s_{+}sin\left(\frac{k\pi\epsilon_{0}}{2}\right)}{k^{2}\pi^{2}\epsilon_{0}}cos\left(\frac{k\pi y}{2}\right)\frac{sinh\left(\frac{k\pi\left(1-x\right)}{2}\right)+sinh\left(\frac{k\pi\left(1+x\right)}{2}\right)}{sinh(k\pi)}$
$-\sum_{k\text{ odd}}\frac{8s_{+}sin\left(\frac{k\pi\epsilon_{0}}{2}\right)}{k^{2}\pi^{2}\epsilon_{0}}cos\left(\frac{k\pi x}{2}\right)\frac{sinh\left(\frac{k\pi\left(1-y\right)}{2}\right)+sinh\left(\frac{k\pi\left(1+y\right)}{2}\right)}{sinh(k\pi)}.$
Proof 4.2.
Since the differential equation and boundary condition of $u_{0}$ is not dependent on $\mathbf{Q}_{0}$, we first solve the differential equation of $u_{0}$. Noting that $u_{0}$ is actually the critical point of the following energy functional,
(69) $E_{0}(u)=\int_{[-1,1]^{2}}|D^{2}u|^{2}\mathrm{d}\mathbf{x},$
with natural boundary conditions. $E_{0}(u)$ is convex on $u$, and thus all the critical points are the global minimiser, i.e. $E_{0}(u_{0})=0$. Consequently, $u_{0}$ satisfies $D_{xx}u_{0}=D_{yy}u_{0}=D_{xy}u_{0}\equiv 0$, which means $u_{0}$ is a linear function,
(70) $u_{0}=k_{1}x+k_{2}y+k_{3},k_{i}\in\mathbb{R},i=1,2,3.$
Given a linear $u_{0}$, the partial differential equation of $\mathbf{Q}_{0}$ is simplified to
(71) $\begin{cases}\Delta\mathbf{Q}_{0}=0,\\ \mathbf{Q}_{0}=\begin{pmatrix}s_{+}L(x)/2&0\\ 0&-s_{+}L(x)/2\end{pmatrix}\text{ on }y=\{1,-1\},\\ \mathbf{Q}_{0}=\begin{pmatrix}-s_{+}L(y)/2&0\\ 0&s_{+}L(y)/2\end{pmatrix}\text{ on }x=\{1,-1\},\end{cases}$
which can be solved by the separation of variables. A standard computation for the WORS profile as in [59] shows the results in (68).
Remark 4.3.
Obviously, the solution of (65) with the boundary condition (66) is not unique because all linear $u_{0}$ and $\mathbf{Q}_{0}$ are the solutions. From the view of energy, the leading term in energy is $|D^{2}u|$ at the $\lambda\rightarrow 0$ limit, and thus it is natural to have a linear or constant $u$ in $\lambda\rightarrow 0$ for vanishing $|D^{2}u|$. The result of $u$ follows the physical intuition that the domain size is too small to accommodate layer structures. One can directly check that the eigenvector of $\mathbf{Q}_{0}$ is either horizontal or vertical, and $\mathbf{Q}_{0}(x,x)=\mathbf{Q}_{0}(x,-x)=0$, which means $\mathbf{Q}_{0}$ has two line defects along the diagonals of square (also see Figure 6).
Now we solve $Q_{1},Q_{2},u_{1},u_{2}$ to see how the state changes with small enough but non-zero $\lambda$. Up to $O(\lambda)$, the governing partial differential equations for $\mathbf{Q}_{1}$ and $u_{1}$ are
(72) $\begin{cases}\Delta^{2}u_{1}=0,\\ \Delta\mathbf{Q}_{1}=0\end{cases}$
with the boundary condition
(73) $\begin{cases}\mathbf{Q}_{1}=0,\text{ on }\partial\Omega,\\ D^{2}u_{1}\cdot\vec{\nu}=0,\left[\nabla\cdot D^{2}u_{1}\right]\cdot\vec{\nu}=0,\text{ on }\partial\Omega.\end{cases}$
which only has the trivial solution, i.e. $Q_{1}\equiv 0$ and linear $u_{1}$. Thus, the leading perturbation is expected to be the second-order term. Up to $O(\lambda^{2})$, the governing partial differential equations for $\mathbf{Q}_{2}$ and $u_{2}$ are
(74) $\begin{cases}\Delta\mathbf{Q}_{2}=A\cdot\mathbf{Q}_{0}+\frac{\text{tr}(\mathbf{Q}_{0}^{2})\mathbf{Q}_{0}}{2}\\ \qquad\qquad+2B_{0}\cdot q^{2}/s_{+}\cdot\left(u_{0}\cdot D^{2}u_{2}-\frac{tr(u_{0}\cdot D^{2}u_{2})}{2}\mathbf{I}_{2}\right)+2B_{0}\cdot q^{4}\cdot\frac{\mathbf{Q}_{0}}{s_{+}^{2}}u_{0}^{2},\\ \Delta^{2}u_{2}=-\nabla\cdot\left(\nabla\cdot\left(q^{2}\left(\frac{\mathbf{Q}_{0}}{s_{+}}+\frac{\mathrm{I}_{2}}{2}\right)u_{0}\right)\right),\end{cases}$
with the boundary condition
(75) $\begin{cases}\mathbf{Q}_{2}=0,\text{ on }\partial\Omega,\\ \left(D^{2}u_{2}+q^{2}\left(\frac{\mathbf{Q}_{0}}{s_{+}}+\frac{I_{2}}{2}\right)u_{0}\right)\cdot\vec{\nu}=0,\left[\nabla\cdot\left(D^{2}u_{2}+q^{2}\left(\frac{\mathbf{Q}_{0}}{s_{+}}+\frac{\mathrm{I}_{2}}{2}\right)u_{0}\right)\right]\cdot\vec{\nu}=0,\text{ on }\partial\Omega.\end{cases}$
The differential equation for $\mathbf{Q}_{2}$ is easy to solve by using the finite difference method, but the differential equation with boundary conditions for $u_{2}$ is difficult to solve for its complex boundary condition, which is a mixture with second and third derivative. Fortunately, the solution of (74) with the boundary condition (75) is actually the critical point for the following energy functional
(76) $\tilde{E}(u_{2})=\int_{[-1,1]^{2}}\left|D^{2}u_{2}+q^{2}\left(\frac{\mathbf{Q}_{0}}{s_{+}}+\frac{\mathbf{I}_{2}}{2}\right)u_{0}\right|^{2}\mathrm{d}\mathbf{x},$
without any boundary anchoring. By minimizing the above energy, we can numerically calculate $u_{2}$, which has oscillation along the director of WORS, as shown in Figure 6(a). At $\lambda^{2}=0.01$, the density distribution, $u$, is no longer a linear function and tends to have a layer structure.
Figure 6: (a) From the left to the right are the profiles of $\mathbf{Q}_{0}=Q_{WORS},u_{0}\equiv 1$, $\mathbf{Q}_{2}$ and $u_{2}$ solved from (74), and $\mathbf{Q}_{0}+0.01\mathbf{Q}_{2},u_{0}+0.01u_{2}$ which is a approximation of the solution at $\lambda^{2}=0.01$. (b) The solution at $\lambda^{2}$=0.01. (c) The plot of difference between $\mathbf{Q}_{0}+0.01\mathbf{Q}_{2},u_{0}+0.01u_{2}$ and the solution ($\mathbf{Q}_{\lambda=0.1},u_{\lambda=0.1}$) at $\lambda^{2}=0.01$, i.e. $\|\mathbf{Q}_{0}+0.01\mathbf{Q}_{2}-\mathbf{Q}_{\lambda=0.1}\|$ and $u_{0}+0.01u_{2}-u_{\lambda=0.1}$. The parameters are $a=-5$, $c=5$, $B_{0}=10^{-3}$, $q=2\pi$, and $A=-0.8359$.
4.3.3 Modest domain size
In this section, we numerically study the confined smectic configuration with modest $\lambda$, which is the complement of $\lambda\rightarrow 0$ and $\lambda\rightarrow\infty$ problems. Unless otherwise specified in the figure caption, the default parameter values are as follows: $a=-5$, $c=5$, $B_{0}=10^{-3}$, $q=2\pi$ (corresponding to a molecular length of approximately $10^{-7}$ m), and $A=-0.8359$ (which is calculated from the parameters in [19, 57, 18]). As the domain size increases from $\lambda^{2}=1$ to $\lambda^{2}=30$, three stable smectic states are shown in Figure 7. These states are the minimisers of (56) and have the lowest energy in our numerical calculations. They have analogous $\mathbf{Q}$ profiles as nematic states, WORS with two line defects on diagonals, BD with two line defects near two opposed edges, and D with no interior defects and directors along one diagonal of the square [57, 18]. The corresponding $u$ profiles have layer normal along the director of $\mathbf{Q}$ profiles. The BD-like and D-like smectic states can be observed in experiments in [16]. The WORS-like state on an extremely small confinement is hard to achieve practically.
Figure 7: The most stable states with various $\lambda^{2}$. From small $\lambda$ to large $\lambda$, the director exhibits WORS, BD, D profile. The colour bar is the same as Figure 4.
It is noticeable that the BD-like state, which is unstable in nematics, becomes stable in smectics. To further explore the effects of positional order on orientational order, we track the branches with small and large $\lambda$ as temperature decreases. In Figure 8(a), for small $\lambda^{2}=4.38$, at high temperatures, the stable state is nematic WORS (where $u\equiv 0$). As the temperature decreases, the central vertical layer structure forms gradually and separates the cross-line defects into two distinct line defects. We speculate that the stability of the BD-like smectic state is enhanced by the positional order profile $u$ to avoid more dislocations in the WORS-like smectic state. In Figure 8(b), for large domain $\lambda^{2}=30$, the nematic D state crystallizes into the smectic D-like state, which indicates the memory of the director in the N-S phase transition.
Figure 8: N-S phase transition as decreasing temperature with square confinement. (a) is the phase transition from WORS-type nematic to BD-type smectic with decreasing temperature at $\lambda^{2}=4.38$, and the rescaled temperature-dependent parameters are $a=1,-0.2,-2,-5$, $A=-0.4286,-0.5916,-0.7544,-0.8359$ from the left to right. (b) is the phase transition from D-type nematic to D-type smectic with decreasing temperature at $\lambda^{2}=30$ and the same rescaled temperature-dependent parameters in (a). The colour bar is the same as Figure 4.
In the nematic phase, when the domain size is large enough, we can find both D and R to be (meta)stable. There are two D states for which the director aligns along the diagonal, and four R states for which the director is rotated by $\pi$ radians between two opposite edges [57, 18]. The profiles of D and R are unique in nematics by taking symmetry into account. However, for smectics, we are able to find multiple (meta)stable D-like and R-like states with subtle differences on $u$ (see Figure 9(a)). This could suggest a frustrated energy landscape, implying the existence of numerous similar minima that differ slightly [60]. By using the saddle dynamics [61], we search the transition pathway between R1 and R2 via an index-1 transition state R3, in Figure 9(b). In such a frustrated energy landscape, it is difficult for an R-like state to break the energy barrier and reach the D-like state with lower energy. One strategy to alleviate the frustration in the energy landscape is to reduce the parameters $a,c,B_{0}$. In Figure 9(c), the energy landscape is smoother as the energy of the $\mathbf{Q}$ field becomes dominant. We find a transition pathway between R-like and D-like states via an index-1 J-like state. This transition pathway is analogous to the nematic counterpart in [18, 33].
Figure 9: (a) R and D type smectic minimum at $\lambda^{2}=30$. The domain enclosed by red lines demonstrates the difference between the two R states and two D states, respectively. (b) A frustated transition pathway with $\lambda^{2}=30,B_{0}=10^{-3},a=-5,c=5$. R1 and R2 are minima, and R3 is an index-1 transition state. R3-R1 (R2) is the slight difference between R3 and R1 (R2). (c) The transition pathway between locally stable R state and more stable D state via index-1 transition state J with $\lambda^{2}=30,B_{0}=10^{-5},a=-0.05,c=0.05$, and the $y$-axis is the scaled energy, $E_{scaled}=e^{E-E(R)}$, for better visualization. The colour bar is the same as Figure 4.
5 Conclusion and discussion
We model the smectic liquid crystal with mLdG free energy. Our study provides an interpretation of the parameters within the mLdG model in terms of physical characteristics, e.g. the coefficient of the quadratic term in the bulk energy of $u$ should be dependent on temperature to model the N-S phase transition, the coefficient of the quartic term in the bulk energy of $u$ determines the amplitude of layer, and the wave number $q$ should be a function of the length of long axis of a rod-like liquid crystal molecule, which allows for a more direct and meaningful comparison with experimental parameters, enhancing the potential of the mLdG model as a predictive tool in liquid crystal research.
More precisely, we first prove the existence and regularity of minimiser. Then, we prove the mLdG energy can model the I-N-S phase transition with respect to temperature, which provides the potential for the application of mLdG in the phase transition. To investigate the influence of the director on the layer structure, we provide preliminary evidence that the mLdG favours the layer structure near the uniaxial area and prefers a non-layered structure near the defect. We also investigate the smectic configuration confined within a square. Our primary findings are as follows: (a) in the limit as $\lambda\rightarrow 0$ or in the case of a very small square, the stable state is the nematic WORS without a layer structure, due to the square being too small to accommodate a layer structure; (b) in the limit as $\lambda\rightarrow\infty$ or in the case of a very large square, there will be an increasing number of smectic layers, given that the height of a layer is constant; (c) for a finite but non-zero $\lambda$ (the experimental setting), the smectic favors the WORS or BD profiles with a small square, but prefers to bend the layer to a D profile with a large square, which is in agreement with experimental results in [16]. We find multiple (meta)stable states without interior defects and the transition pathways between them, which demonstrates the frustration in the energy landscape. The N-S phase transitions in square confinement show the influence of the layer structure on the defect and the memory of the director in the N-S phase transition.
The results of this paper suggest numerous open questions. By using colloidal silica rods and leveraging their significant density difference with the dispersing solvent, isotropic, nematic, and smectic phases can be confined within a single chamber [16]. This phenomenon could be modelled using the 3D mLdG model, as it allows for phase transitions within the third dimension by setting $z$-dependent $a$ and $A$ in (1), i.e. $a=a(z)$ and $A=A(z)$. Additionally, within three-dimensional confinement, we expect to find a greater variety of smectic configurations, as the nematic confined within a 3D cuboid presents a richer phase diagram compared to a 2D square [43]. In [16], the smectic configurations confined between two spherical shells exhibit a chevron structure, which is an interface between two disoriented domains at the outer sphere. It would be intriguing to numerically determine the mechanism behind such a chevron structure. In future research, we plan to explore the smectic configuration confined between two spherical shells.
Acknowledgements
This work was partially supported by Royal Society Newton Advanced Fellowship NAF/R1/180178 awarded to AM and LZ. LZ is supported by the National Natural Science Foundation of China (grants 12225102, T2321001, 12050002, 12288101). BS would like to thank the University of Strathclyde for its support and hospitality when work on this paper was undertaken. AM is supported by the University of Strathclyde New Professors Fund, University of Strathclyde GEF-ISP grant, the Humboldt Foundation, a Leverhulme Research Project Grant RPG-2021-401, and the INI network grant supported by EP/R014604/1. YH is supported by a Leverhulme Research Project Grant RPG-2021-401.
Appendix: Numerical method
In this section, we describe the numerical methods used to compute the (meta)stable state, i.e. minimiser, and the transition state, i.e. index-1 saddle of the given energy. Typically, a stable state can be easily found by the gradient descent method using a proper initial guess, and a transition state can be found by the saddle dynamic [61] using a minimiser as an initial guess. For the time discretization, we employ a Barzilai-Borwein (BB) step size [62], which exhibits a superlinear convergence rate. In the confinement problem, we give the tangential boundary condition of $\mathbf{Q}$, and it is convenient to use finite difference methods for spatial discretization with mesh size $\delta x$. However, special attention must be given to the time discretization. The BB step size could potentially introduce instability due to the fourth-order differential operator within our framework. To address this instability issue, we only specify our numerical scheme within the gradient flow for the reason that the saddle dynamics can be discretized in the same way. The discretization of the gradient flow of (1) is,
(77) $\frac{\mathbf{Q}_{n+1}-\mathbf{Q}_{n}}{\Delta t_{n}}=$ $-K\Delta_{\delta x}\mathbf{Q}_{n}-A\cdot\mathbf{Q}_{n}-C\cdot tr(\mathbf{Q}_{n}^{2})\mathbf{Q}_{n}$
$-2B_{0}\cdot q^{2}/s_{+}\cdot\left(u_{n}\cdot D^{2}_{\delta x}u_{n}-\frac{tr(u\cdot D^{2}_{\delta x}u_{n})}{2}\mathbf{I}_{2}\right)-2\cdot B_{0}\cdot q^{4}\cdot\frac{\mathbf{Q}_{n}}{s_{+}^{2}}u_{n}^{2},$
$\frac{u_{n+1}-u_{n}}{\Delta t_{n}}=$ $-2B_{0}\Delta_{\delta x}^{2}u_{n+1}-au_{n}-cu_{n}^{3}-4B_{0}\cdot D^{2}_{\delta x}u_{n}:\left(q^{2}\cdot\left(\frac{\mathbf{Q}_{n}}{s_{+}}+\frac{\mathbf{I}_{2}}{2}\right)\right)$
$-2B_{0}\cdot\nabla_{\delta x}\cdot\left(\nabla_{\delta x}\cdot\left(q^{2}\cdot\left(\frac{\mathbf{Q}_{n}}{s_{+}}+\frac{\mathbf{I}_{2}}{2}\right)u_{n}\right)\right)$
$-4B_{0}\cdot\left|q^{2}\cdot\left(\frac{\mathbf{Q}_{n}}{s_{+}}+\frac{\mathbf{I}_{2}}{2}\right)\right|^{2}u_{n},$
where $\Delta_{\delta x}^{2},\Delta_{\delta x},\nabla_{\delta x}$, $D^{2}_{\delta x}$ are the discretization of $\Delta^{2},\Delta,\nabla,D^{2}$, and $\Delta t_{n}$ is the BB step size at the $n$-th iteration. In (77), we discretize the fourth-order operator $\Delta^{2}$ implicitly to ensure the stability of the BB step size. Conversely, we discretize the second-order operator $\Delta$ explicitly, as it is compatible with the BB step size in our numerical computations and offers the advantage of saving computational resources.
In Section 3, we study the phase transition with periodic boundary conditions, and we use the spectral method [55] for spatial discretization,
(78) $Q(x)=\sum_{k=-N/2}^{N/2}\tilde{Q}_{k}e^{2\pi ikx/h},Q\in V_{Q},u(x)=\begin{cases}\sum_{k=-N/2}^{N/2}\tilde{u}_{k}e^{2\pi ikx/h},u\in V_{u},\\ \sum_{k=1}^{N+1}\tilde{u}_{k}\sin\left(2k\pi x/h\right),u\in V,\end{cases}$
where $N$ is an even integer. Recall that $V=V_{u}\cap W^{1,2}_{0,\Omega}$, so we use the sine spectral method to discretize $u\in V$. By substituting (78) in (23), we obtain a discretized form of the energy,
(79) $E(\tilde{Q}_{k},\tilde{u}_{k})\approx E(Q,u)$
This results the functional (78) in a function of $2(N+1)$ dimensions, and we directly search the minimum by using the gradient descent method [63] for finite-dimensional functions.